# Amplifying Main Memory-Based Timing Covert and Side Channels using Processing-in-Memory Operations

### v3

@@ 5.3 Covert-Channel Attack Throughput Analysis
Fig. 8 shows the throughput achieved by different covert-channel attacks when one sender and one receiver thread run concurrently across systems with increasing LLC sizes. We make five key observations. First, IMPACT-PnM and IMPACT-PuM achieve significantly higher throughput than all other attack vectors. IMPACT-PnM and IMPACT-PuM achieve communication throughputs of 12.87 Mb/s and 14.16 Mb/s, respectively, irrespective of the cache size. They support up to 4.91$\times$ and 5.41$\times$ of the throughput achieved by the state-of-the-art main memory-based attack, DRAMA-clflush. Second, IMPACT-PuM achieves 10% higher throughput than that of IMPACT-PnM since the sender can issue all RowClone requests in parallel, whereas in IMPACT-PnM, the sender has to send the PiM instructions in a sequential manner. Third, as the LLC size increases, the throughput of DRAMA-eviction, DRAMA-clflush, and Streamline decreases significantly. This is because the larger the LLC size, the higher the latency it takes to access the LLC. Fourth, the DMA engine attack achieves a throughput of 5.27 Mb/s irrespective of the cache size but is almost 2.4$\times$ slower than IMPACT-PnM due to its OS-related overheads (e.g., context switch overhead, OS instructions). Fifth, the PnM-OffChip attack achieves a maximum throughput of 12.64 Mb/s. As the LLC size increases, the throughput is reduced to 10.64 Mb/s because the off-chip predictor decides to cache more data when the LLC is large. We conclude that IMPACT achieves significantly higher throughput than the existing state-of-the-art main memory-, cache- and DMA engine-based covert-channel attacks.
Figure 8: Throughput across increasing LLC sizes.
Breakdown of Throughput Improvements. Fig. 9 shows the time (in CPU cycles) it takes for the sender to send the message and for the receiver to read the message in IMPACT-PnM and IMPACT-PuM. We observe that in IMPACT-PuM, the sender can send the message 14$\times$ faster than the sender in IMPACT-PuM since it can issue all RowClone requests in parallel. In IMPACT-PnM, even though the sender is significantly slower than the sender of IMPACT-PuM, the sender and the receiver overlap, at the cost of synchronization, to hide the latency of the sender. This is why IMPACT-PnM achieves only 10% lower throughput than IMPACT-PuM.
@@ 5.4 Side-Channel Attack Throughput Analysis
Fig. 10 shows the throughput and the error rate of the side-channel attack against read mapping (RM) across different numbers of DRAM banks in the left and right y-axis, respectively. We measure the throughput of the attack based on the correct guesses and calculate the error rate based on the number of incorrect guesses. The exact end-to-end accuracy of identifying the properties of the sample genome depends on the chaining mechanism and the imputation algorithm [110, 111] used by the attacker, and it is out of the scope of this work.
We make two key observations. First, with a PiM-enabled DRAM with 1024 banks, the attacker can leak information about the sample genome at a throughput of 7.57 Mb/s and an error rate of $<5$%. Second, as the number of banks increases to 8196, the throughput of the attack decreases to 2.56 Mb/s and the error rate increases to $<15$%. This is because the attacker has to probe more banks, which makes the attack more prone to noise and reduces its bandwidth.
However, at the same time, the leaked information provides more exact information about the sample genome since the number of hash table entries per bank decreases, reducing the number of potential locations of the reads in the reference genome per bank. For example, in a PiM-enabled DRAM device with 1024 banks, the victim activates a row with 16 hash table entries. Assuming the attacker identifies the activated bank correctly, they determine that one of the 16 hash table entries was targeted by the victim. In contrast, in a PiM-enabled device with 2048 banks, the victim activates a row with 8 hash table entries. When the attacker identifies the activated bank correctly, they determine that one of the 8 hash table entries was targeted by the victim, which leaks more precise information on the sample genome.
Figure 10: Leakage throughput and error rate of the RM side-channel attack across different numbers of DRAM banks.
@@ 6 Defenses against IMPACT
This section discusses the effectiveness of three defense mechanisms that eliminates the row buffer-based timing channel.
@@
Memory Partitioning (MPR). Bank-level memory partitioning enforces that each DRAM bank can only be accessed by one process that it is allocated to. In IMPACT, this limits the sender’s ability to transmit information across different DRAM banks and the receiver’s ability to observe contention created by the sender. However, this approach has three major drawbacks. First, it can lead to limiting the number of concurrently running applications due to the fixed number of DRAM banks available. Second, it can cause memory underutilization due to allocating DRAM bank-sized memory to applications. Third, it disables data sharing, which can significantly increase data duplication and memory underutilization, and lead to frequent swapping from the disk when the memory space is full.
@@
Closed Row Policy (CRP). Enforcing a memory controller policy that closes a DRAM row after each operation can be an effective defense against IMPACT. By closing the DRAM row after each operation, the memory controller ensures that subsequent memory accesses cannot access the same row without requiring a row activation. Closed row policy disrupts the timing patterns exploited by the attacker in IMPACT. However, it also leads to significant performance overheads, even for workloads with low data locality, as every DRAM access leads to a row buffer miss.
@@
Constant-Time DRAM Accesses (CTD). Another defense mechanism is to enforce a constant-time DRAM access (CTD) policy. When a memory access is issued, the memory controller waits for a fixed amount of time before returning the data to the processor. This fixed amount of time is equal to the worst-case DRAM access latency. By enforcing CTD, the memory controller ensures that all memory accesses take the same amount of time, regardless of the memory location. Thus, obfuscates the timing patterns exploited by the attacker in IMPACT. However, this leads to significant performance overheads as every DRAM access has the worst-case latency.
Performance Evaluation. We evaluate the performance of CRP and CTD compared to the baseline open-row policy for 5 multi-programmed graph applications which (i) share the input graph (i.e., access to the same DRAM bank) and (ii) experience a mix of low and high data locality (i.e., the number of last-level cache misses per kilo instructions). We use a 2-core system where each core runs a different instance of the same application (e.g., core 0 runs an instance of PageRank and core 1 the same instance of PageRank using the same input). We evaluate five workloads from GraphBIG benchmark suite [120]: Betweenness Centrality (BC; Data set size: 7.8 GB; LLC MPKI: 0.57), Breadth-First search (BFS; 7.6 GB; 38.59), Connected components (CC; 7.9 GB; 45.2), Triangle counting (TC; 8.2 GB; 5.08), PageRank (PR; 7.0 GB; 1.86).
Fig. 11 shows the performance overhead of CRP and CTD compared to the open-row policy. We make two key observations. First, CTD leads to the highest performance degradation, 26% on average, compared to the open row policy. Second, CRP leads to an average performance degradation of 15% compared to the open row policy but leads to low performance overheads in workloads that do not benefit from the open row policy such as TC, CC and BFS. We conclude that fundamental defense mechanisms that eliminate the timing channel lead to high performance overheads, and there is a dire need for high-performance defense mechanisms.
Figure 11: Performance overhead of Closed-Row policy and Constant-Time policy compared to Open-Row policy.
@@ 7 Related Work
To the best of our knowledge, this is the first work that evaluates the timing attack vulnerabilities in PiM systems.
@@
Security Implications of PiM Systems. A set of prior works [121, 122, 123] discuss specific security vulnerabilities in PiM systems. Arafin et al. [121] survey and theoretically analyze security vulnerabilities, focusing on five specific attacks. In this work, we quantitatively and qualitatively analyze and evaluate main memory-based covert and side channels in detail. Wang et al. [123] extracts neural network architectural information from an RRAM accelerator using power side channels. Ensan et al. [122] proposes an attack on RRAM-based PiM architectures using power/timing side channels. In contrast, we demonstrate how to establish high-throughput timing covert and side channels using PiM operations.
@@
Other Main Memory-based Timing Attack Defenses. We already extensively discuss defenses that fundamentally eliminate the timing channel (§6) and defenses based on cache management methods that are inapplicable. Another approach is to restrict the fine-grained timers’ usage to prevent attackers from measure memory requests’ latencies, which is applied in some modern processors (e.g., Apple M1 cores [124]). However, these timers are used in many user space applications (e.g., performance and power consumption monitoring, multithreading, and synchronization and locking). Restricting access to these timers can disable these applications or degrade their performance. DAGguise [125] is a defense mechanism against memory timing side-channel attacks that obfuscates an application’s memory access pattern. DAGguise does not mitigate row buffer-based timing channels and requires a closed row policy. Trusted execution environments (TEEs) (e.g., Intel SGX [126, 127]) can isolate applications from the rest of the system. However, prior works already showcase multiple side-channel attacks on TEEs leveraging the main memory [128], and other structures [129, 130, 131, 128, 132, 133, 134, 135, 136, 137].
@@
Enhancing System Security using PiM. Several prior works enhance the security by leveraging PiM. SecNDP [138] proposes an encryption and verification mechanism for untrusted PiM devices. It only focuses on protecting confidentiality and does not consider or analyze potential timing channels. Nelson et al. [139] propose offloading security critical tasks to near-memory compute units to avoid storing sensitive information in caches and eliminate cache-based side-channel attacks. In this work, we demonstrate and evaluate the implications of PiM on main memory-based timing attacks. InvisiMem [140] and ObfusMem [141] propose leveraging 3D-stacked memory to design secure processors that provide oblivious RAM equivalent guarantees.
@@ 8 Conclusion
We introduce IMPACT, a set of high-throughput main memory timing attacks that leverage the direct memory access enabled by PiM architectures. IMPACT exploits the shared DRAM row buffer to 1) establish covert channels to communicate with other processes at high throughput and 2) leak security-critical information of victim processes based on their memory accesses. IMPACT covert channels leverage parallelism and eliminate cache bypassing steps, resulting in increased throughput compared to existing main memory-based covert channels. IMPACT side-channel attack targets a genomics application and leaks the private characteristics of a user’s sample genome by leveraging PiM operations. We evaluate and discuss three fundamental defense mechanisms against IMPACT.

### v4

@@ 6.2 Covert Channel Throughput Analysis
We make four key observations. First, IMPACT-PnM and IMPACT-PuM achieve significantly higher throughput than all other attack vectors. IMPACT-PnM and IMPACT-PuM provide communication throughputs of 8.2 Mb/s and 14.8 Mb/s, respectively, irrespective of the cache size. They support up to 3.6$\times$ and 6.5$\times$ of the throughput achieved by the state-of-the-art main memory-based attack, DRAMA-clflush. Second, IMPACT-PuM provides 82% higher throughput than IMPACT-PnM. This is because the sender in IMPACT-PuM can transmit a message with lower latency using fewer operations (§4.2). Third, as the LLC size increases, the throughputs of DRAMA-eviction and DRAMA-clflush decrease significantly due to the increasing LLC access latency. Fourth, the DMA engine attack provides a throughput of 0.81 Mb/s irrespective of the LLC size, which is 10.1$\times$ slower than IMPACT-PnM due to its OS-related overheads (e.g., context switch overhead, OS instructions). We conclude that IMPACT-PnM and IMPACT-PuM achieve significantly higher throughput than the existing state-of-the-art main memory-based covert-channel attacks.
Figure 9: Leakage throughput of IMPACT-PnM, IMPACT-PuM, and state-of-the-art main memory-based covert channel attacks.
Breakdown of Throughput Improvements. We analyze the throughput breakdown of IMPACT-PnM and IMPACT-PuM by calculating the time spent on sender and receiver routines to transmit a fixed-length message. Fig. 10 shows the time (in CPU cycles) it takes for the sender to send the message and for the receiver to decode it in IMPACT-PnM and IMPACT-PuM. We make two key observations. First, in IMPACT-PuM, the sender routine takes 11.1$\times$ less time than the sender in IMPACT-PnM. This is because IMPACT-PuM can issue only one RowClone request to transmit N bits, where N is the number of banks. In contrast, IMPACT-PnM transmits the same message with potentially multiple PEIs. Second, the receiver routines of IMPACT-PnM and IMPACT-PuM spend similar amounts of time decoding the message as they both need to use distinct PiM operations to detect which DRAM banks were accessed by the sender. Based on these observations, we conclude that IMPACT-PuM has a higher throughput than IMPACT-PnM due to the differences in the sender routines.
@@ 6.3 Side-Channel Attack Throughput Analysis
We evaluate the side-channel attack against genomic read mapping (RM) (§4.3) in a PiM-enabled system [67], sweeping the number of DRAM banks that store the hash table used in the seeding step. Fig. 11 shows the throughput and the error rate of the side-channel attack in the primary y-axis and secondary y-axis, respectively. The x-axis shows the total number of DRAM banks in the system. We measure the throughput of the attack based on the correct guesses of the hash table entries accessed and calculate the error rate based on the number of incorrect guesses. [^3] The exact end-to-end accuracy of identifying the properties of the query genome depends on the chaining mechanism and the imputation algorithm [160, 161] used by the attacker, and it is out of the scope of this work.
Figure 11: Leakage throughput and error rate of the genomic read mapping side-channel attack.
We make two key observations. First, the attacker can leak information about the query genome at a throughput of 7.57 Mb/s and an error rate of $<5$% in a system with 1024 banks. Second, as the number of banks increases, the throughput of the attack decreases to 2.56 Mb/s and the error rate increases to $<15$%. This is because the attacker has to probe more banks, making the attack more prone to noise and reducing its bandwidth.
However, at the same time, the leaked information provides more exact information about the query genome since the number of hash table entries per bank decreases, reducing the number of potential locations of the reads in the reference genome per bank. For example, in a PiM-enabled DRAM device with 1024 banks, the victim activates a row with 16 hash table entries. Assuming the attacker identifies the activated bank correctly, they determine that one of the 16 hash table entries was targeted by the victim. In contrast, in a PiM-enabled device with 2048 banks, the victim activates a row with 8 hash table entries. When the attacker identifies the activated bank correctly, they determine that one of the 8 hash table entries was targeted by the victim, which leaks more precise information on the query genome.
Based on these observations, we conclude that an attacker can successfully leak query genome characteristics in a PiM-enabled system by leveraging PiM operations.
@@ 7 Defenses Against IMPACT
We demonstrate high-throughput attack vectors that can emerge with the adoption of PiM architectures. Our results provide insights into new side and covert channel attack vectors that PiM systems might face, and show that these attack vectors should be considered while designing such PiM systems. In this section, we discuss the effectiveness of four defense mechanisms that mitigate the row buffer-based timing channel we introduce.
7.1 Memory Partitioning (MPR)
@@
Bank-level memory partitioning (similarly to [168, 169]) allocates each DRAM bank to only one process and enforces that only that process accesses the DRAM bank. This way, none of the processes can observe latency differences due to other processes’ memory accesses. Thus, the attacker processes cannot build communication channels or leak information. However, this approach has three major drawbacks. First, it limits the number of applications that can run concurrently due to the fixed number of DRAM banks available in the system. Second, it can cause memory underutilization due to allocating DRAM bank-sized memory to applications. Third, it disables data sharing across processes, which can significantly increase data duplication and memory underutilization, and lead to frequent swapping from the disk when the memory space is full.
7.2 Closed Row Policy (CRP)
@@
Enforcing a memory controller policy that closes a DRAM row after each operation, a closed row policy (similar to as described in [155, 170, 171]), can be an effective defense against IMPACT. By closing the DRAM row after each operation, the memory controller ensures that subsequent memory accesses cannot access the same row without requiring a row activation. The closed row policy disrupts the timing patterns exploited by IMPACT. However, it also leads to significant performance overheads [172, 170], even for workloads with low data locality, as every DRAM access leads to a row buffer miss.
7.3 Constant-Time DRAM Accesses (CTD)
@@
Another defense mechanism is to enforce a constant-time DRAM access (CTD) policy. When a memory access is issued, the memory controller waits for a fixed amount of time before returning the data to the processor. This fixed amount of time is equal to the worst-case DRAM access latency. By enforcing CTD, the memory controller ensures that all memory accesses take the same amount of time, which eliminates the timing differences exploited by IMPACT. However, this leads to significant performance overheads (§7.4) as every DRAM access has the worst-case latency.
7.4 Reducing the Attack Throughput with Adaptive Constant-Time DRAM (ACT)
Another approach is to reduce the leakage throughput (without completely getting rid of the covert-channel attack) while incuring lower performance overheads than existing defenses. To understand the effectiveness of such a solution, we implement a memory controller-based countermeasure that adaptively enforces constant-time memory access latency per DRAM bank based on the row buffer contention observed. We call this countermeasure Adaptive Constant-Time (ACT). ACT counts the row buffer conflicts for each DRAM bank during an epoch and decides which latency policy (i.e., constant-time or default latency) to employ in the next one. It switches to the constant-time policy for a bank when there are more row buffer conflicts than a predetermined threshold in the last epoch. This way, in the next epoch, all memory requests being served by the bank will exhibit the same latency (i.e., the worst case latency), and the attacker always observes interference. While employing constant-time latency, ACT still keeps track of the row buffer conflicts. It switches back to the default policy only when the number of conflicts is lower than the threshold in the last epoch. This can reduce the performance overheads caused by CTD. ACT does not fundamentally solve the row buffer-based timing channel, and an attacker can still mount the attack by transmitting messages only when ACT is employing default latency and then waiting while ACT employs constant-time latency. However, this limits the usable epochs during execution and reduces the attack throughput.
We implement ACT and evaluate its impact on system performance and the attack throughput. We evaluate three different configurations an epoch is set to 1000ns: (i) ACT-Aggressive: employing constant-time latency for the next 4000 epochs after the 1st conflict occurs in the bank, (ii) ACT-Mild: employing constant-time latency for the next 2 epochs after the 1st conflict occurs in the bank, (iii) ACT-Conservative: employing constant-time latency for the next 2 epochs after 5 conflicts occur in the bank. To analyze ACT’s impact on system performance, we evaluate four workloads from the GraphBIG benchmark suite [173]: Betweenness Centrality, Breadth-First search (BFS), Connected components (CC), Triangle counting (TC) and, XSBench (XS) [174]. To analyze ACT’s impact on attack throughput, we calculate the reduction in IMPACT-PnM’s throughput. Figure 12 shows the performance comparison of ACT-Conservative, ACT-Mild, ACT-Aggressive, and CTD normalized to a system with no IMPACT defense.
Figure 12: Performance comparison of Constant-Time policy and Adaptive Constant-Time defense policies over the baseline system with no IMPACT defense.
We make two key observations. First, ACT-Aggressive causes similar performance overhead as CTD, while reducing the throughput of IMPACT-PnM by 72% on average (not shown in the figure), in contrast to CTD, which prevents the attack completely. Second, the two less aggressive variants of ACT (ACT-Mild and ACT-Conservative) lead to only 10% performance overhead. However, they cannot reduce the throughput of IMPACT-PnM. This is because the attack utilizes all available banks during a batch of transmission, which increases the idle time between each transmission that uses a specific DRAM bank (i.e., the attack uses the same bank only after a time period where it issues requests to the other banks). This requires significantly high delay times for the mitigation to reduce the throughput. Our findings highlight the necessity for a new approach in defending against main-memory timing attacks.
8 Discussion
8.1 Other Potential PiM-Based Attack Vectors
In this work, we show that emerging PiM architectures can enable a set of high-throughput main memory-based timing attacks based on the DRAM row buffer. Adopting PiM architectures can lead to other attack vectors or exacerbate existing ones, requiring further exploration and analysis. One such example is the impact of PiM on read disturbance vulnerability [175, 172]. A prior PiM proposal in 2021 [82] states that the proposal may lead to increased vulnerability to RowHammer [175] attacks. A new work in 2025, PuDHammer [176], studies the read disturbance effects of Processing-using-DRAM (PuD) operations and shows that these operations greatly increase read disturbance vulnerabilities of DRAM modules (up to 158.6$\times$). Thus, it is critical to employ robust and scalable RowHammer [175] and RowPress [172] solutions to prevent read disturbance bitflips in PiM-enabled systems.
Other potential attack vectors include 1) thermal attacks that leverage PiM operations to increase the temperature in DRAM to make DRAM unavailable and restrict its service time, and degrade reliability, and 2) timing attacks that leverage coarse-grained function offloading. Exploring these directions requires special attention and research, which we leave for future work.
8.2 Applicability of IMPACT and its Defenses
to Complex PiM Architectures
Our observations apply to other PiM architectures as long as they satisfy the key properties that enable the attack: 1) allowing co-locating data from different applications in a DRAM bank, and 2) enabling subsequent accesses from different applications targeting the same DRAM bank. Complex PnM architectures (e.g., systems like Tesseract [177, 178, 179] that employ sophisticated prefetching and network functionality) can degrade the attack throughput by introducing additional noise.
The performance overhead of an IMPACT defense heavily depends on how the defense mechanism interacts with the underlying system. The defense mechanisms we discuss in §7 are not specifically tailored for a given underlying system (e.g., constant-time policy). The new defense mechanism that we introduce in §7.4 enables the constant-time policy only for the bank under contention, reducing the performance overhead, but does not provide strong security guarantees.
8.3 Restricting Access to PIM Operations
A key contribution of our study is to show that unrestricted access to PiM operations enables potentially harmful high-throughput side and covert channel attack vectors. None of the current PiM proposals we are aware of discuss restricting access to PiM operations to ensure security. We hope and believe that our study will guide system designers in balancing the performance, flexibility, and security trade-offs while implementing effective access control mechanisms to PiM.
8.4 Applicability to Future DRAM Devices
The row buffer timing channel would persist in future DRAM-based systems, enabling IMPACT. With newer generations, we observe that DRAM modules consist of increasingly more banks, which increases IMPACT’s covert channel throughput as the attack would benefit from more parallelism. Another change is the introduction of RowHammer mitigations, PRAC [180, 181, 182] and RFM [183], in recent DDR specifications that block DRAM accesses when necessary to prevent RowHammer bitflips. These mitigations incur at least 350ns (at most 1400ns) latency per preventive action. This is significantly higher than the row buffer conflict latency and can be filtered out by the receiver.
We believe that increasing DRAM’s complexity will increase attack vectors as manufacturers try to combat the challenges of technology scaling. For example, recent preprints show that PRAC and RFM introduce timing channels that can be exploited for side-channel [184] and covert-channel [184, 185] attacks.
@@ 9 Related Work
To our knowledge, this is the first work that evaluates timing attack vulnerabilities in PiM systems.
9.1 Security Implications of PiM Systems
@@
A set of prior works [186, 187, 188] discuss specific security vulnerabilities in PiM systems. Arafin et al. [186] survey and theoretically analyze security vulnerabilities, focusing on five specific attacks. In this work, we quantitatively and qualitatively analyze and evaluate main memory-based covert and side channels in detail. Wang et al. [188] extract neural network architectural information from an RRAM accelerator using power side channels. Ensan et al. [187] propose an attack on RRAM-based PiM architectures using power/timing side channels. In contrast, we demonstrate how to establish high-throughput timing covert and side channels using PiM operations.
9.2 Other Timing Attacks
We already discuss main memory-based timing attacks in §2.3. In addition, prior works showcase multiple cache-based timing side and covert-channel attacks (e.g., [189, 190, 191, 192, 193, 194, 195, 196, 197]). Among these works, Streamline [192] has the highest throughput as it avoids explicit cache flushes and utilizes a large transmission set to evict cache lines implicitly by thrashing the cache. IMPACT achieves a leakage throughput within the same order of magnitude of Streamline and does not require a transmission set larger than the cache size.
9.3 Other Main Memory Timing Attack Defenses
@@
We already extensively discuss defenses that fundamentally eliminate the timing channel (§7). Another approach is to restrict the high-resolution timers’ usage to prevent attackers from measuring memory requests’ latencies, which is applied in some modern processors (e.g., Apple M1 cores [198]). However, these timers are used in many userspace applications (e.g., performance and power consumption monitoring, multithreading, and synchronization and locking). Restricting access to these timers can disable these applications or degrade their performance. DAGguise [199] is a defense mechanism against memory timing side-channel attacks that obfuscates an application’s memory access pattern. DAGguise does not mitigate row buffer-based timing channels and requires a closed row policy. Trusted execution environments (TEEs) (e.g., Intel SGX [200, 201]) can isolate applications from the rest of the system. Prior works [202, 203, 204, 205, 206, 207, 208, 209, 210, 211] already demonstrate multiple side-channel attacks on TEEs leveraging main memory [205], cache hierarchy [202, 203, 204, 205, 206], page table entries [207, 208, 209, 205], and branch predictors [210, 211].
9.4 Enhancing System Security using PiM
@@
Several prior works enhance system security by leveraging PiM. SecNDP [212] proposes an encryption and verification mechanism for untrusted PiM devices. It only focuses on protecting confidentiality and does not consider or analyze potential timing channels. Nelson et al. [213] propose offloading security-critical tasks to PnM units to avoid storing sensitive information in caches and eliminate cache-based side-channel attacks. In this work, we demonstrate and evaluate the implications of PiM on main memory-based timing attacks. InvisiMem [214] and ObfusMem [215] propose leveraging 3D-stacked memory to design secure processors that provide oblivious RAM equivalent guarantees.
@@ 10 Conclusion
We introduce IMPACT, a set of high-throughput main memory timing attacks that leverage the direct memory access enabled by Processing-in-Memory (PiM) architectures. IMPACT exploits the shared DRAM row buffer to 1) establish covert channels to communicate with other processes at high throughput and 2) leak security-critical information of victim processes based on their memory accesses. IMPACT covert channels leverage parallelism and eliminate cache bypassing steps, resulting in increased throughput compared to existing main memory-based covert channels. IMPACT side-channel attack targets a genomics application and leaks the private information with a low error rate. We evaluate and discuss four potential defense mechanisms against IMPACT and eventually conclude that mitigating IMPACT incurs high performance overhead and more research is needed to find low-overhead solutions.
Acknowledgments
We thank the anonymous reviewers of USENIX Security’24, ISCA’24, MICRO’24, ASPLOS’25, and DSN’25 for their constructive feedback. We thank the SAFARI Research Group members for providing a stimulating intellectual environment. We acknowledge the generous gifts from our industrial partners, including Google, Huawei, Intel, and Microsoft. This work, along with our broader work in Processing-in-Memory and memory systems, is supported in part by the Semiconductor Research Corporation (SRC), the ETH Future Computing Laboratory (EFCL), AI Chip Center for Emerging Smart Systems (ACCESS), sponsored by InnoHK funding, Hong Kong SAR, European Union’s Horizon programme for research and innovation [101047160 - BioPIM], a Google Security and Privacy Research Award, and the Microsoft Swiss Joint Research Center.
Appendix A Artifact Appendix
A.1 Abstract
This artifact provides the source code and scripts to reproduce the experiments and results presented in our DSN 2025 paper. It includes implementations of various covert channel attacks (PNM, DRAMA-clflush, DRAMA-evict, PUM, DMA) and a defense mechanism, along with scripts to automate the setup, execution, and result analysis.
A.2 Artifact Checklist (Meta-information)
  • Program: C++ programs, Python scripts, shell scripts.
  • Compilation: GNU Make.
  • Run-time environment: Linux (tested on Ubuntu 20.04 and 22.04), Python 3.
  • Execution: Bash scripts, Python scripts.
  • Metrics: Timing results, performance metrics of attacks and defenses.
  • Output: Figures in PDF format and related data in plaintext files.
  • How much disk space required (approximately)?: 20GB.
  • How much time is needed to prepare workflow (approximately)?: $\sim 1$ hour.
  • How much time is needed to complete experiments (approximately)?: 2-4 hours.
  • Publicly available?: Yes
  • Archived (provide DOI)?: Yes, DOI: 10.5281/zenodo.15116851
A.3 Description
A.3.1 How to Access
The source code and scripts can be downloaded from Zenodo (https://zenodo.org/records/15116851) and GitHub (https://github.com/CMU-SAFARI/IMPACT).
A.3.2 Hardware dependencies
  • x86-64 system.
  • At least 8GB of RAM (128GB recommended for parallel defense experiments).
  • $>64$ cores are suggested for parallel defense experiments.
A.3.3 Software dependencies
  • Linux Operating System (tested on Ubuntu 20.04 and 22.04).
  • Bash.
  • Python (version 3.12.9)
  • GNU Make (version 4.3).
  • wget, tar.
  • GCC (11.4.0), g++ (11.4.0).
  • zlib1g-dev.
A.3.4 Data sets
The defense mechanism experiments may require trace files. The script includes commented-out lines to download example traces. You may need to download these or provide your own traces.
A.4 Installation
A.4.1 Dependency Installation
To install all dependencies, run the dependency installation script: install_dependencies.sh.
{mdframed}
[backgroundcolor=darkgray, linecolor=black, linewidth=0.5pt] \$ cd IMPACT
\$ sh install_dependencies.sh
\$ bash
\$ conda activate impact
A.4.2 Conda Environment Setup
If you would like to install Miniconda manually, follow the steps below.
  1. 1.
  2. 2.
    Run the Miniconda installer.
  3. 3.
    Open a new terminal.
A.5 Experiment Workflow
To run all experiments, run the main script: artifact.sh. This script 1) builds the attacks and the simulator and 2) runs all experiments required to reproduce the key results.
{mdframed}
[backgroundcolor=darkgray, linecolor=black, linewidth=0.5pt] \$ sh artifact.sh
The artifact.sh script will execute the following steps.
A.5.1 Building the Attacks and Simulator
  1. 1.
    The script will navigate to the impact/covert_channel_attack/ directory.
  2. 2.
    It will clean and build the attack code using make clean and make all.
  3. 3.
    It will navigate to the simulator/sniper directory.
  4. 4.
    It will clean and build the simulator using make distclean and make -j\$(nproc).
A.5.2 Running Examples and Attacks
A.5.3 Basic Example Run
The script includes a commented-out line to run a basic example using the Sniper simulator. Uncomment the line.
A.5.4 Running Attacks for Figure 9
The script executes the following attacks:
  • PNM: python3 \$home_dir/impact/covert_channel_attack/script/run_pnm.py \$home_dir 16 \$home_dir/results_attacks/pnm
  • DRAMA (clflush): python3 \$home_dir/impact/covert_channel_attack/script/run_drama.py \$home_dir 16 \$home_dir/results_attacks/drama
  • DRAMA (eviction sets): python3 \$home_dir/impact/covert_channel_attack/script/run_drama_ev.py \$home_dir 16 \$home_dir/results_attacks/drama_ev
  • PUM: python3 \$home_dir/impact/covert_channel_attack/script/run_pum.py \$home_dir 16 \$home_dir/results_attacks/pum
  • DMA: python3 \$home_dir/impact/covert_channel_attack/script/run_dma.py \$home_dir 16 \$home_dir/results_attacks/dma
A.5.5 Running Attacks for Figure 8 and 10 (PNM POC)
python3 \$home_dir/impact/covert_channel_attack/script/run_pnm_poc.py \$home_dir 16 \$home_dir/results_attacks/pnm_poc
A.5.6 Running Attacks for Figure 8 and 10 (PUM POC)
python3 \$home_dir/impact/covert_channel_attack/script/run_pum_poc.py \$home_dir 16 \$home_dir/results_attacks/pum_poc
A.5.7 Running Defense Mechanism Experiments for Figure 12
  • Download traces (if needed).
  • Create the jobfile: python3 \$home_dir/scripts/create_jobfile_impact.py \$home_dir/simulator/sniper \$home_dir/scripts/defenses.jobfile
  • Execute the jobfile: sh \$home_dir/scripts/defenses.jobfile
A.6 Plotting the Results
Once the experiments are completed, the reader can plot the figures using the plotting scripts we provide.
To reproduce the figures, 1) install plotting dependencies:
{mdframed}
[backgroundcolor=darkgray, linecolor=black, linewidth=0.5pt] \$ pip install numpy matplotlib pandas
2) run the plotting script: plot_figures.sh
{mdframed}
[backgroundcolor=darkgray, linecolor=black, linewidth=0.5pt] \$ sh plot_figures.sh
This script will create Figures 8, 9, 10, and 12 under the figures/ directory.
A.7 Evaluation & Expected Results
A.7.1 Post-Execution Steps
Review Figures 8, 9, 10, and 12 in the figures/ directories.
Note that the attack codes are compiled during the setup, and the results presented in Figures 8, 9, and 10 might vary slightly based on your system configurations (e.g., compiler version).
A.7.2 Troubleshooting
Refer to the troubleshooting section in the README.md for potential issues and solutions.
A.8 Methodology
Submission, reviewing and badging methodology: